OpenStudy (anonymous):

solve -1 = -2v + 2/3 for v.

OpenStudy (anonymous):

was thtat worded like that in the problem

OpenStudy (anonymous):

yahh

OpenStudy (anonymous):

on your sheet or computer

OpenStudy (anonymous):

i mean no

OpenStudy (anonymous):

can you copy and paste it??or rite it how the problem was ??

OpenStudy (anonymous):

mmkay

OpenStudy (mathstudent55):

Is this it? Solve for v \(-1 = -2v + \dfrac{2}{3} \)

OpenStudy (anonymous):

yeah @mathstudent55

OpenStudy (anonymous):

-2/3 then correct @mathstudent55

OpenStudy (mathstudent55):

First, subtract 2/3 from both sides. Then, divide both sides by -2. gtg. try it, i'll be back in 20 min.

OpenStudy (anonymous):

ugh can i have the answer lol? @dirtydan667

OpenStudy (anonymous):

thts against the policy of openstudy i could get banned

OpenStudy (anonymous):

and i dont want tht

OpenStudy (anonymous):

really? oh its soo hard though

OpenStudy (anonymous):

ill show youtep by step with the awnser just one sec

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

for some stupid reason i cant get it

OpenStudy (anonymous):

can you figure the answer out? like explain it to me with the answer in it lol

OpenStudy (mathstudent55):

\(-1 = -2v + \dfrac{2}{3} \) We want v by itself, so first, subtract 2/3 from both sides. \(-1 - \dfrac{2}{3} = -2v + \dfrac{2}{3} - \dfrac{2}{3}\) \(-\dfrac{3}{3} - \dfrac{2}{3} = -2v + \dfrac{2}{3} - \dfrac{2}{3}\) Do the subtractions on both sides. What do you get?

OpenStudy (anonymous):

i dont know help please! :(

OpenStudy (anonymous):

oh 1/3 = -2v right? @mathstudent55

OpenStudy (mathstudent55):

\(-\dfrac{3}{3} - \dfrac{2}{3} = -2v + \dfrac{2}{3} - \dfrac{2}{3} \) This is the next step. \( -\dfrac{5}{3} = -2v \) \(-2v = -\dfrac{5}{3} \) Now you need to divide both sides by -2, which is the same as multiplying both sides by -1/2.

OpenStudy (anonymous):

i got 5 thats not right help! @mathstudent55

OpenStudy (mathstudent55):

\( -2v = -\dfrac{5}{3} \) \( -\dfrac{1}{2} \times ( \dfrac{-2v}{1}) = -\dfrac{1}{2} \times ( -\dfrac{5}{3} ) \) \( v = \dfrac{5}{6} \)

OpenStudy (anonymous):

THANKS

OpenStudy (mathstudent55):

You're welcome.